OpenStudy (anonymous):

how do i solve a question like : -2=6 over 7 p?? Thanks :) P.S will give medal for best answer <3

OpenStudy (jiteshmeghwal9):

\[-2={6 \over 7p}\]

OpenStudy (anonymous):

so what do i do ??? to solve it :)

OpenStudy (jiteshmeghwal9):

\[7p \times -2=6\]

OpenStudy (jiteshmeghwal9):

\[-14p=6\]

OpenStudy (jiteshmeghwal9):

\[p={6 \over -14p}\]

OpenStudy (jiteshmeghwal9):

\[p=\frac{3 \cancel6}{-7 \cancel{-14}}=\frac{3}{-7}ans.\]

OpenStudy (jiteshmeghwal9):

\[\frac{-3}{7} < 3\]

OpenStudy (jiteshmeghwal9):

gt it @fLuFfYzOiNgObOots ??

OpenStudy (anonymous):

no sorry :)

OpenStudy (jiteshmeghwal9):

wht is ur doubt ??

OpenStudy (anonymous):

on my answer sheet is says that it will be -2 and one third

OpenStudy (jiteshmeghwal9):

\[-2 \frac{1}{3}\]is this ur answer

OpenStudy (anonymous):

no the books answer

OpenStudy (jiteshmeghwal9):

if it is then look